# Quantitative entropy estimates for 2D stochastic vortex model on the whole space under moderate interactions

### 05 Mar 2026 v3

@@ 3.1 Time evolution of energy
$\left\|\rho_{t}-\rho^{N}_{t}\right\|_{2}^{2}$ $\leq\left\|\rho_{0}-\rho^{N}_{0}\right\|_{2}^{2}+\widetilde{I}^{1}_{t}-\widetilde{M}_{t}^{N}-2\int_{0}^{t}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds$
@@
3.2 Estimates for $\widetilde{I}_{t}$: the nonlinear terms
Now by by subtracting and adding the term $\rho_{s}^{N}K\ast\rho_{s}^{N}$, along with integration by parts, we have
$2\int_{\mathbb{R}^{2}}(\rho_{s}-\rho_{s}^{N})\left[\nabla\cdot\left[\rho_{s}K\ast\rho_{s}\right]-\nabla\cdot\left[\left\langle S_{s}^{N},V^{N}(x-\cdot)K\ast\rho_{s}^{N}(\cdot)\right\rangle\right]\right]\,dx$
$=-2\int_{\mathbb{R}^{2}}\nabla(\rho_{s}-\rho_{s}^{N})\left[\rho_{s}K\ast\rho_{s}-\left\langle S_{s}^{N},V^{N}(x-\cdot)K\ast\rho_{s}^{N}(\cdot)\right\rangle\right]\,dx$
$=-2\int_{\mathbb{R}^{2}}\nabla(\rho_{s}-\rho_{s}^{N})\left[\rho_{s}K\ast\rho_{s}-\rho_{s}^{N}K\ast\rho^{N}_{s}\right]\,dx$
@@ 3.2 Estimates for $\widetilde{I}_{t}$: the nonlinear terms
In addition, since $\nabla\cdot K=0$, by by subtracting and adding the terms $\rho_{s}$ and $\rho_{s}^{N}K\ast\rho_{s}$, we find
@@
By Young’s inequality, we observe that
@@
$\|\rho_{s}-\rho_{s}^{N}\|_{4}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}\|\rho_{s}\|_{2}$ $\lesssim\|\rho_{s}-\rho_{s}^{N}\|_{2}^{1/2}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{1/2}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}\|\rho_{s}\|_{2}$
$\leq C_{\epsilon}\|\rho_{s}\|_{2}^{4}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}.$ (88)
@@
Now we will take care about $\widetilde{I}^{1,2}$ in (85). First, taking into account (7), with analogous computations such as in section 2 of [40], we derive
@@
with $\widetilde{\theta}_{2}\doteq\beta(2-3\gamma)$ and $\widetilde{\theta}_{3}\doteq\frac{\beta\gamma^{2}}{d}$, for $\gamma\in(1/2,2/3)$.
@@
$\lesssim 2\|\rho_{s}^{N}\|_{1}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})|\left(N^{-\widetilde{\theta}_{2}}+N^{-\widetilde{\theta}_{3}}\rho_{s}^{N}\right)$
$+2\|\rho_{s}^{N}\|_{4}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})|\left(N^{-\widetilde{\theta}_{2}}+N^{-\widetilde{\theta}_{3}}\rho_{s}^{N}\right)$
@@
Now, by by subtracting and adding $\rho_{s}$, we note that
@@
$+N^{-\widetilde{\theta}_{3}}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})||\rho_{s}-\rho_{s}^{N}|$
$+N^{-\widetilde{\theta}_{3}}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})||\rho_{s}|$
@@
$\widetilde{I}^{1,2,2}$ $\leq N^{-\widetilde{\theta}_{2}}\|\rho_{s}^{N}\|_{4}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{1}+N^{-\widetilde{\theta}_{3}}\|\rho_{s}^{N}\|_{4}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})|\rho_{s}^{N}$
@@
3.3 End of the proof: application of Gronwall Lemma
First, with the same computations as in [40], for all $\delta\in(0,\widetilde{\theta}_{4})$ there exists a random variable $\widetilde{A}_{0}$, finite almost surely, such that
@@ 3.3 End of the proof: application of Gronwall Lemma
Then by (98) and (99) in (84), for $\epsilon<<1$, we obtain
$\left\|\rho_{t}-\rho^{N}_{t}\right\|_{2}^{2}+\int_{0}^{t}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds$ $\leq\left\|\rho_{0}-\rho^{N}_{0}\right\|_{2}^{2}$
$+C_{\epsilon}\int_{0}^{t}\|\rho_{s}\|_{2}^{2}\|\rho_{s}-\rho_{s}^{N}\|_{1}^{2}\,ds$
@@
$\overset{}{+}N^{-\widetilde{\theta}_{2}}\int_{0}^{t}\left(\|\nabla\rho_{s}\|_{1}+\|\nabla\rho_{s}^{N}\|_{1}\right)\,ds$
$\overset{}{+}2N^{-\widetilde{\theta}_{2}}\int_{0}^{t}\left(\|\nabla\rho_{s}\|_{1}^{2}+\|\nabla\rho_{s}^{N}\|_{1}^{2}\right)\,ds$
$+N^{-\widetilde{\theta}_{3}}C_{\epsilon}\int_{0}^{t}\|\rho_{s}\|_{2}^{2}\,ds+N^{-\widetilde{\theta}_{1}}\int_{0}^{t}\|\rho_{s}\|_{4}^{2}\,ds$
@@
Now by (18) $\|\rho\|_{T,2},\|\rho\|_{T,4}\leq C_{0}$, almost surely, and thus for a universal constant $C>0$ in (100) along side (60), (2.3) and Csiszár-Kullback-Pinsker inequality (107), we derive
$\left\|\rho_{t}-\rho^{N}_{t}\right\|_{2}^{2}+\int_{0}^{t}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds$
$\hskip 100.37498pt\leq\left\|\rho_{0}-\rho^{N}_{0}\right\|_{2}^{2}$
@@
$\hskip 100.37498pt\overset{}{+}CN^{-\widetilde{\theta}_{2}}\int_{0}^{t}\|\nabla\rho_{s}\|_{1}\,ds+CN^{-\widetilde{\theta}_{2}}\int_{0}^{t}\|\nabla\rho_{s}\|_{1}^{2}\,ds$
$\hskip 100.37498pt\overset{}{+}CN^{-\widetilde{\theta}_{1}}$
@@
$\hskip 100.37498pt+CN^{-\widetilde{\theta}_{3}}+CN^{-\widetilde{\theta}_{2}}$
@@
Finally, by adding (65) and (101), we deduce for $\epsilon<<1$ and $N>\left(\frac{4}{\epsilon}\right)^{1/\widetilde{\theta}_{1}}$
@@
$C_{N}\doteq CN^{-\widetilde{\theta}}\|\nabla\rho\|_{T,1}+CN^{-\widetilde{\theta}}\|\nabla\rho\|_{T,1}^{2}+CN^{-\widetilde{\theta}}+CA_{0}N^{-\widetilde{\theta}}+C\widetilde{A}_{0}N^{-\widetilde{\theta}}$,
@@
$\widetilde{\theta}\doteq\min{\left(\theta_{1},\theta_{2},\widetilde{\theta}_{1},\widetilde{\theta}_{2},\widetilde{\theta}_{3},\widetilde{\theta}_{4}\right)}-\delta$
@@
It follows by (15), (23), (24) and (81), (99) in (102), that there exist $T>0$ and $\widetilde{\Lambda}\in\mathcal{F}$, $\mathbb{P}(\widetilde{\Lambda})=1$, with the following property: for all $\omega\in\widetilde{\Lambda}$, there exists $N_{0}=N_{0}(\omega)$ such that, if $N\geq N_{0}$ we have $N^{-\widetilde{\theta}}+C_{N}+TC<\exp{(-CT)}(2CT)^{-1/2}$ and

### 13 Apr 2026 v4

@@ 2.2 Time evolution of energy
$\left\|\rho_{t}^{N}-\rho_{t}\right\|_{2}^{2}$ $\leq\left\|\rho_{0}^{N}-\rho_{0}\right\|_{2}^{2}+\widetilde{I}^{1}_{t}-\widetilde{M}_{t}^{N}-2\int_{0}^{t}\left\|\nabla(\rho_{s}^{N}-\rho_{s})\right\|_{2}^{2}\,ds$
@@
2.2.1 Estimates for $\widetilde{I}_{t}$: the nonlinearity
We now address the nonlinearity in the evolution equation of energy. The main novelty is identifying, within the energy estimates, the Fisher information of the particle system alongside Ladyzhenskaya inequality, to address the nonlinearity.
By subtracting and adding the term $\rho_{s}^{N}K\ast\rho_{s}^{N}$, we have
$\widetilde{I}^{1}_{t}=-2\int_{\mathbb{R}^{2}}\nabla(\rho_{s}-\rho_{s}^{N})\left[\rho_{s}K\ast\rho_{s}-\rho_{s}^{N}K\ast\rho^{N}_{s}\right]\,dx$
@@ 2.2.1 Estimates for $\widetilde{I}_{t}$: the nonlinearity
In addition, since $\nabla\cdot K=0$, by subtracting and adding the term $\rho_{s}$, we find
@@
By Young inequality, we observe that
@@
$\|\rho_{s}-\rho_{s}^{N}\|_{4}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}\|\rho_{s}\|_{2}$
$\hskip 80.29999pt\lesssim\|\rho_{s}-\rho_{s}^{N}\|_{2}^{1/2}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{1/2}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}\|\rho_{s}\|_{2}$
$\hskip 80.29999pt\leq C_{\epsilon}\|\rho_{s}\|_{2}^{4}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}.$ (63)
@@
Now we will take care about $\widetilde{I}^{1,2}$ in (60).
First, by decay of $V$, given by (6), we have, for $|x-\cdot|\geq N^{-\beta\gamma/d}$
$V^{N}(x-\cdot)|x-\cdot|^{\gamma}$ $=N^{\beta}V(N^{\beta/d}(x-\cdot))|x-\cdot|^{\gamma}$
[…]
$\lesssim N^{-3+\beta}\left|(x-\cdot)\right|^{-3d/\beta}$
$\lesssim N^{-\widetilde{\theta}_{2}},$ (65)
which implies that
@@
with $\widetilde{\theta}_{2}\doteq 3-\beta-3\gamma$ and $\widetilde{\theta}_{3}\doteq\frac{\beta\gamma^{2}}{d}$, for $\gamma\in(0,1/2)$.
@@
$\lesssim 2\|\rho_{s}^{N}\|_{1}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})|\left(N^{-\widetilde{\theta}_{2}}+N^{-\widetilde{\theta}_{3}}\rho_{s}^{N}\right)\,dx$
$+2\|\rho_{s}^{N}\|_{4}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})|\left(N^{-\widetilde{\theta}_{2}}+N^{-\widetilde{\theta}_{3}}\rho_{s}^{N}\right)\,dx$
@@
Now, by subtracting and adding $\rho_{s}$ along side Young inequality, we note that
@@
$+N^{-\widetilde{\theta}_{3}}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})||\rho_{s}-\rho_{s}^{N}|\,dx$
$+N^{-\widetilde{\theta}_{3}}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})||\rho_{s}|\,dx$
@@
$\widetilde{I}^{1,2,2}$ $\leq N^{-\widetilde{\theta}_{2}}\|\rho_{s}^{N}\|_{4}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{1}+N^{-\widetilde{\theta}_{3}}\|\rho_{s}^{N}\|_{4}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})|\rho_{s}^{N}\,dx$
@@
2.3 Combining entropy and energy
First, with the same computations as in [25], for all $\delta\in(0,\theta_{3})$ there exists a random variable $A_{0}$, finite almost surely, such that
$\sup_{t\in[0,T]}|M_{t}^{N}|\leq A_{0}N^{-\theta_{3}+\delta},$ (75)
with $\theta_{3}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)$.
In addition, with the same computations as in [36], for all $\delta\in(0,\widetilde{\theta}_{4})$ there exists a random variable $\widetilde{A}_{0}$, finite almost surely, such that
@@ 2.3 Combining entropy and energy
[…]
$\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})-\mathcal{H}(\rho_{0}^{N}|\rho_{0})$ $\leq C\int_{0}^{t\wedge\tau^{N}}\left(\mathcal{H}(\rho_{s}^{N}|\rho_{s})+\|\rho^{N}_{s}-\rho_{s}\|_{2}^{2}\right)\,ds$
$\overset{}{+}(-1+2\epsilon)\int_{0}^{t\wedge\tau^{N}}\left(\mathcal{I}(\rho_{s}^{N}|\rho_{s})+\|\nabla(\rho_{s}^{N}-\rho_{s})\|_{2}^{2}\right)\,ds$
$+CN^{-\theta_{1}}+CN^{-\theta_{2}}+A_{0}N^{-\theta_{3}+\delta},$ (77)
for a universal constant $C>0$.
Analogously, by (74) and (76) in (59), we obtain
$\left\|\rho_{t}^{N}-\rho_{t}\right\|_{2}^{2}-\left\|\rho_{0}^{N}-\rho_{0}\right\|_{2}^{2}$ $\leq C_{\epsilon}\int_{0}^{t}\|\rho_{s}\|_{2}^{2}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds$
@@
$\overset{}{+}2N^{-\widetilde{\theta}_{2}}\int_{0}^{t}\left(\|\nabla\rho_{s}\|_{1}^{2}+\|\nabla\rho_{s}^{N}\|_{1}^{2}+T\right)\,ds$
$+N^{-\widetilde{\theta}_{3}}C_{\epsilon}\int_{0}^{t}\|\rho_{s}\|_{2}^{2}\,ds+N^{-\widetilde{\theta}_{2}}\int_{0}^{t}\|\rho_{s}\|_{4}^{2}\,ds$
@@
Now, since $\rho_{s}^{N}\in\mathcal{P}\left(\mathbb{R}^{d}\right)$, $\mathbb{P}$-a.s. and then by Holder’s inequality we obtain
Also, from Leibniz’s rule we deduce,
$\overset{}{\leq}2\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+C_{\eta}+2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\left|\nabla\ln\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx,$ (80)
where in the last step we are using the following: for $\eta>2C_{1}(8T+C_{3})+1$
$\int_{\mathbb{R}^{d}}\frac{C_{3}}{1+t}\exp{\left(-\frac{|x-X_{t}|^{2}}{8t+C_{3}}\right)}\exp{\left(\frac{C_{1}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx<\infty$
which implies by Donsker-Varadhan inequality (130),
[…]
Hence since (17) implies $\|\rho\|_{L^{\infty}L^{2}},\|\rho\|_{L^{\infty}L^{4}}\leq C_{0}$, almost surely, by (78), (79) and (81), there exists a universal constant $C>0$ in (78), we derive
$\left\|\rho_{t}^{N}-\rho_{t}\right\|_{2}^{2}+\int_{0}^{t}\left\|\nabla(\rho_{s}^{N}-\rho_{s})\right\|_{2}^{2}\,ds$
$\hskip 100.37498pt\leq\left\|\rho_{0}^{N}-\rho_{0}\right\|_{2}^{2}$
@@
$\hskip 100.37498pt\overset{}{+}CN^{-\widetilde{\theta}_{2}}\int_{0}^{t}\|\nabla\rho_{s}\|_{1}^{2}\,ds+TCN^{-\widetilde{\theta}_{2}}$
@@
$\hskip 100.37498pt\overset{}{+}CN^{-\widetilde{\theta}_{1}}+CN^{-\widetilde{\theta}_{2}}+CN^{-\widetilde{\theta}_{3}}$
@@
It follows that, by adding (77) and (82), for $\epsilon<<1$ and $N>>1$, we deduce
@@
$C_{N}\doteq CN^{\widetilde{\theta}}\|\nabla\rho\|_{L^{\infty}L^{1}}^{2}+CN^{\widetilde{\theta}}+CA_{0}N^{\widetilde{\theta}}+C\widetilde{A}_{0}N^{\widetilde{\theta}}$,
@@
$\widetilde{\theta}\doteq\min{\left(\theta_{1},\theta_{2},\theta_{3},\widetilde{\theta}_{1},\widetilde{\theta}_{2},\widetilde{\theta}_{3},\widetilde{\theta}_{4}\right)}-\delta$
@@
2.4 End of the proof: Removal the stopping time
$=\left\{\exists i;\sup_{t\in[0,T]}|X_{t}^{i,N}|\geq N^{\beta}\right\}$
$\subset L_{1}^{N}\cup L_{2}^{N}\cup L_{3}^{N},$ (85)
with
and
We note that, with the same computations like those led (34), we deduce
We now take care about the $L_{2}^{N}$. First, we observe that, in view of Definition 5.1 in Chapter 2 of [24],
$\Omega=\prod_{j=1}^{d}\Omega^{j},\,\,\,\mathcal{F}=\bigotimes_{j=1}^{d}\mathcal{F}^{j},\,\,\,\mathbb{P}=\bigotimes_{j=1}^{d}\mathbb{P}^{j}$ (90)
and then
$W_{t}^{i,N}(\omega)=(W_{t}^{i,N,j}(\omega_{j}))_{j=1}^{d}$ (91)
for $W^{i,N,j}$ a standard, one-dimensional Brownian motion on $(\Omega^{j},\mathcal{F}^{j},\mathbb{P}^{j})$, $j\in\{1,...,d\}$.
Therefore, we note that $W_{t}^{i,N}=\int_{0}^{t}dW_{s}^{i,N}$, and since
$|W_{t}^{i,N}|^{2}=\sum_{j=1}^{d}(W_{t}^{i,N,j})^{2},$
if $\sup_{t\leq T}|W_{t}^{i,N}|\geq a$, then there exists $j\in\{1,...,d\}$, such that $\sup_{t\leq T}|W_{t}^{i,N,j}|\geq\frac{a}{\sqrt{d}}.$
Hence,
$\left\{\sup_{t\leq T}|W_{t}^{i,N}|\geq a\right\}\subset\bigcup_{j=1}^{d}\left[\Omega^{1}\times...\times\widetilde{\Omega}^{j}\times...\times\Omega^{d}\right]$ (92)
with
$\widetilde{\Omega}^{j}\doteq\left\{\sup_{t\leq T}|W_{t}^{i,N,j}|\geq\frac{a}{\sqrt{d}}\right\}$
and then, we obtain
[…]
$\leq\left(\sqrt{\frac{T}{2\pi}}\frac{4dN}{\frac{C_{K_{0},V}N^{\beta}}{\sqrt{d}}}\right)\exp{\left(-\left(\frac{C_{K_{0},V}N^{\beta}}{\sqrt{d}}\right)^{2}/(2T)\right)},$ (93)
where in the second inequality, we are using (8.3)’ in Chapter 2 of [24], page 96.
Now we will derive an analogous estimate for $L_{3}^{N}$. First, we observe that, if $\sigma=0$ we are done, thus suppose $\sigma\neq 0$. Define,
$Y_{t}\doteq\int_{0}^{t}\sigma_{s}\,dB_{s},$ (94)
which implies
$Y_{t}^{j}=\sum_{k=1}^{d}\underbrace{\int_{0}^{t}\sigma_{s}^{j,k}\,dB_{s}^{k}}_{\doteq Y_{s}^{j,k}}.$ (95)
Since $\sigma:[0,T]\to\mathbb{R}^{d}\times\mathbb{R}^{d}$ is bounded, for $j\in\{1,...,d\}$, the function defined by $t\mapsto\langle Y^{j,k}\rangle_{t}$ is continuous and increasing.
Hence, by Theorem 4.6, in Chapter 3 of [24], for all $j\in\{1,...,d\}$, we get
$\sup_{t\leq T}|Y_{t}^{j}|$ $\leq\sum_{k=1}^{d}\sup_{t\leq T}\left|\int_{0}^{t}\sigma_{s}^{j,k}\,dB_{s}^{k}\right|\leq\sum_{k=1}^{d}\sup_{t\leq T}\left|\widetilde{B}^{j,k}_{\langle Y^{j,k}\rangle_{t}}\right|$
$\leq\sum_{k=1}^{d}\sup_{t\leq\|\sigma\|_{\infty}^{2}T}|\widetilde{B}^{j,k}_{t}|,$ (96)
where $(\widetilde{B}^{j,k}_{t})_{t\geq 0}$ is the time-change process, which in particular, is a one dimensional standard Brownian motion. In addition, since
$|Y_{t}|^{2}=\sum_{j=1}^{d}|Y_{t}^{j}|^{2}$
if $\sup_{t\leq T}|Y_{t}|\geq a$, then there exists $j\in\{1,...,d\}$, such that $\sup_{t\leq T}|Y_{t}^{j}|\geq\frac{a}{\sqrt{d}}.$ Therefore, by (90) and for all $a>0$, we find
$\left\{\sup_{t\leq T}|Y_{t}|\geq a\right\}$ $\subset\bigcup_{j=1}^{d}\left\{\sup_{t\leq T}|Y_{t}^{j}|\geq\frac{a}{\sqrt{d}}\right\}$
[…]
$\overset{}{\subset}\bigcup_{j=1}^{d}\bigcup_{k=1}^{d}\left[\Omega^{1}\times...\times\widetilde{\Omega}^{j,k}\times...\times\Omega^{d}\right],$ (97)
with
$\widetilde{\Omega}^{j,k}\doteq\left\{\sup_{t\leq\|\sigma\|_{\infty}^{2}T}|\widetilde{B}^{j,k}_{t}|\geq\frac{a}{d^{3/2}}\right\}.$ (98)
It follows that, by (96) and (97)
[…]
$\leq\left(\sqrt{\frac{\|\sigma\|_{\infty}^{2}T}{2\pi}}\frac{4d^{2}N}{\frac{C_{K_{0},V}N^{\beta}}{d^{3/2}}}\right)\exp{\left(-\left(\frac{C_{K_{0},V}N^{\beta}}{d^{3/2}}\right)^{2}/(2\|\sigma\|_{\infty}^{2}T)\right)},$ (99)
where in the second inequality, we are using (8.3)’ in Chapter 2 of [24], page 96.
Hence, by (89), (93) and (99) in (85), we find
$\sum_{N\in\mathbb{N}}\mathbb{P}\left(\tau^{N}<T\right)<\infty$ (100)
and then by Borel-Cantelli Lemma, we arrive at
Finally, by (14), (23), (19), (20), (37), (75), (76) and (101) in (83) , for $T<\min{\left(T_{1};(1/\|\nabla V\|_{1}\|K_{0}\|_{\infty})\right)}$ for all $\omega\in\Lambda$, with
$\Lambda\doteq\Lambda_{0}\cap\left(\bigcup_{M=1}^{\infty}\bigcap_{N=M}^{\infty}\{\tau^{N}=T\}\right)\cap\left\{A_{0}<\infty\right\},$ (102)
there exists $N_{0}(\omega)$ such that for $N\geq N_{0}(\omega)$, we have $\tau^{N}(\omega)=T$,
$\mathcal{H}(\rho_{0}^{N}|\rho_{0})+\|\rho_{0}-\rho_{0}^{N}\|_{2}^{2}+C_{N}<\exp{(-CT)}(2CT)^{-1/2}$
and